Corrections etc

Note the two changes in the syllabus introduced last year.

(1) Secret sharing introduced.

(2) You now need to be able to give an informal account of quantum cryptography.

Note also that the Shannon’s theorems in the syllabus are those given in the notes. Some earlier lecturers went deeper into entropy than this lecturer and there are one or two earlier exam questions which go beyond what will be done in the present course.

The abreviation GVS should be replaced by GSV where it occurs. (I think page 27 only.)